OpenStudy (anonymous):

The girls in Karen’s family own a total of 46 T-shirts. The oldest, Heidi, has 18 T-shirts. Karen has six times as many T-shirts as her little sister, June. How many T-shirts does Karen own? The girls in Karen’s family own a total of 46 T-shirts. The oldest, Heidi, has 18 T-shirts. Karen has six times as many T-shirts as her little sister, June. How many T-shirts does Karen own? @Mathematics

OpenStudy (anonymous):

Another translation question. Read the problem carefully, and you will get the following equation:\[46=18+K+\frac{K}{6}\]Solve for K, and you are golden.

OpenStudy (anonymous):

\[28=\frac{7K}{6} \implies K=24\]

OpenStudy (anonymous):

Seems Karen has about twenty-four t-shirts.